### v1

# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

Abstract
Large Language Model (LLM)-based multi-agent systems (MAS) demonstrate remarkable potential for scientific discovery. Existing approaches, however, often automate scientific discovery using predefined workflows that lack rationality constraints. This often leads to aimless hypothesizing and a failure to consistently link hypotheses with evidence, thereby hindering systematic uncertainty reduction. Overcoming these limitations fundamentally requires systematic uncertainty reduction. We introduce PiFlow, an information-theoretical framework, treating automated scientific discovery as a structured uncertainty reduction problem guided by principles (e.g., scientific laws). In evaluations across three distinct scientific domains – discovering nanomaterial structures, bio-molecules, and superconductor candidates with targeted properties – our method significantly improves discovery efficiency, reflected by a 73.55% increase in the Area Under the Curve (AUC) of property values versus exploration steps, and enhances solution quality by 94.06% compared to a vanilla agent system. Overall, PiFlow serves as a Plug-and-Play method, establishing a novel paradigm shift in highly efficient automated scientific discovery, paving the way for more robust and accelerated AI-driven research. Code is publicly available at our GitHub.
1 Introduction
Large Language Model (LLM)-based Multi-Agent Systems (MAS) have significantly impacted automated scientific discovery Minaee2024LargeLM ; Wang2023ASO ; Zhang2024ACS across a wide range of fundamental fields, including chemistry Liu2024DrugAgentAA ; Ghafarollahi2024AtomAgentsAD ; Yang2024MOOSEChemLL ; Inoue2024DrugAgentED , biology Xiao2024CellAgentAL ; Nagarajan2025IANAI ; Averly2025LIDDIALI ; Ghafarollahi2024ProtAgentsPD , physics Jaiswal2024ImprovingPR , and material science Takahara2025AcceleratedIM ; Ghafarollahi2025AutomatingAD ; Ansari2024dZinerRI ; Wan2024FromTT ; Zhang2024LargeLM .
Although proficient in executing experiments within predefined workflows Lu2024TheAS ; lai2025prim , these systems often generate hypotheses that lack clear direction, leading to the uncertainty establishing clear links between hypotheses and their supporting or refuting evidence AI4Science2023TheIO ; Zhou2024HypothesisGW ; Baek2024ResearchAgentIR ; Schmidgall2025AgentLU ; Prabhakar2025OmniScienceAD ; Xie2023DARWINSD . Such a disconnect indicates inefficient exploration. Moreover, many of these approaches are tailored for specific tasks, often relying on meticulous prompt engineering that heavily incorporates domain knowledge (as detailed in Appendix A). Consequently, their ability to adapt to new scientific domains is often limited without substantial modifications Zhang2025ScientificLL ; Kumbhar2025HypothesisGF . These issues culminate in three primary challenges: (a) aimless hypothesizing; (b) unmaintained connections between hypotheses and evidence during exploration, hindering systematic validation and (c) limited generalization ability, where systems effective in one scenario (e.g., material science) often require substantial rework to be applicable in others.
To address these limitations, we introduce PiFlow, an information-theoretical framework for structured uncertainty reduction in scientific discovery. Viewing scientific exploration as a game against an unknown and challenging nature, where robust strategies are paramount, PiFlow employs Min-Max optimization: minimizing cumulative regret for exploitation, and maximizing information gain for hypothesis exploration. As a Plug-and-Play module, PiFlow integrates with MAS capable of hypothesizing and experimentation. Inspired by the challenge of navigating vast hypothesis spaces, PiFlow operates by using fundamental scientific principles, which may be initially proposed by or refined using LLMs, to steer the exploration.
The iterative selection of principles progressively reduces uncertainty in hypothesizing and in the interpretation of evidence, dynamically steering exploration by prioritizing those scientific principles that offer the highest instructive value for continued exploration. Figure 1 illustrates PiFlow’s method for assessing and utilizing scientific principles within a drug discovery context.
Figure 1: Illustration of the potential of a scientific principle in drug discovery. PiFlow directs exploration to prioritize hypotheses aligned with high-potential principles (or their variants), thereby iteratively guiding the discovery towards optimal candidate molecules.
This principle-aware approach yields systematic information gain: PiFlow selects high-potential principles and then guides hypothesizing via three actions, i.e., exploring, validating, or refining their scope and formulation. Thus, PiFlow progressively optimizes its guiding scientific principles to effectively steer hypothesizing. Furthermore, leveraging its Min-Max optimization, PiFlow theoretically achieves cumulative regret growth of $\mathcal{O}(\sqrt{T})$ over $T$ exploration steps. This sublinear regret underscores its guaranteed efficiency in navigating complex discovery landscapes. In summary, our contributions are:
  1. (a)
    We propose a novel paradigm for principle-aware scientific discovery, built upon an information-theoretical foundation that offers theoretical convergence guarantees.
  2. (b)
    We develop PiFlow, a Plug-and-Play framework that seamlessly integrates with existing MAS to enable focused exploration and exploitation, thereby enhancing discovery efficiency and operational flexibility.
  3. (c)
    We conduct extensive experiments across three distinct scientific discovery scenarios, demonstrating the broad applicability and significant performance improvements achieved by PiFlow.
2 Related Work
2.1 Language Models for Scientific Discovery
Recently, large language models (LLM) have advanced scientific discovery with automation and rational design (Ren2025TowardsSI, ; Ma2024LLMAS, ). The internal knowledge of LLMs has demonstrated promising capability in focused chemical and material discovery (Yang2024MOOSEChemLL, ; Zhou2024HypothesisGW, ; pu2024leveraging, ; Ghafarollahi2024RapidAA, ). While tool-integrated LLMs like SciAgents (Ghafarollahi2024SciAgentsAS, ), DARWIN (Xie2023DARWINSD, ) and HoneyComb (Zhang2024HoneyCombAF, ) improve domain-specific reasoning and recall of factual insights, they still struggle to integrate physicochemical laws effectively when refining insights for design, risking biased proposals and inefficient exploration due to inherent hallucinations of LLMs (Zhang2023HowLM, ). Human-AI frameworks address this issue by leveraging the knowledge of domain experts (Reddy2024TowardsSD, ; Eythorsson2025TowardAS, ), yet remain limited to the scope of hypothesis generation (Alkan2025ASO, ), leading to insufficient exploration of complex chemical spaces (Luo2025LLM4SRAS, ). Surveys highlight persistent gaps in efficiency and interpretability (Zhang2024ACS, ; Han2024FromGT, ), underscoring the need for principled scientific discovery management beyond automated LLM reasoning (Ramos2024ARO, ).
2.2 Approaches of Multi-Agent Collaboration
Multi-agent systems (MAS) show promise for complex tasks (Lu2024TheAS, ; Ghafarollahi2024SciAgentsAS, ; Ni2023MechAgentsLL, ), yet their application to scientific discovery reveals limitations in current collaboration mechanisms (Tran2025MultiAgentCM, ). Rule-based methods (Zhang2023ExploringCM, ) offer consistency but their predefined rules lack the flexibility to incorporate nuanced scientific principles (e.g., physicochemical laws) or adapt to unexpected findings, hindering dynamic exploration. Role-playing approaches (Tran2025MultiAgentCM, ; He2024LLMBasedMS, ) leverage agent expertise, yet rigid roles can impede adaptation in scientific research (RamirezMedina2025AcceleratingSR, ; Lu2024TheAS, ; Ghafarollahi2024SciAgentsAS, ), and ensuring collective adherence to scientific principles when interpreting experimental insights is challenging. Model-based methods (Xu2023MAgICIO, ; Mu2023RuntimeVO, ; Li2023TheoryOM, ) aim for adaptability by learning from uncertainty, but struggle to build world models that accurately capture complex scientific phenomena and integrate guiding laws (Hao2023ReasoningWL, ), thereby impairing the balance between information perception and strategic, principle-guided reasoning.
Consequently, existing MAS paradigms often lack a dedicated awareness and systematic application of scientific principles during hypothesis generation and refinement (Gridach2025agentic, ; Luo2025LLM4SRAS, ; Reddy2024TowardsSD, ; Su2024ManyHA, ). This highlights a critical need for an explicitly principle-aware multi-agent collaboration framework. Our work addresses this gap, proposing a method where the collaborative discovery process is robustly guided by scientific principles to achieve more efficient and reliable outcomes.
3 Methodology
3.1 Overview
We propose a principle-aware Multi-Agent System (MAS) designed to enhance scientific discovery via focused hypothesizing and structured exploration of hypothesis-evidence connections. Figure 2 illustrates the architecture. Its core comprises a Hypothesis-Validation loop that iteratively generates and tests hypotheses. PiFlow guides this loop by analyzing accumulated principle-outcome data. Strategic insights from PiFlow are relayed through a Planner agent ($\mathcal{A}_{P}$) to the Hypothesis Agent within the loop. Subsequent sections will elaborate on PiFlow’s specific architecture and its information-theoretical underpinnings.
3.2 Architecture
Our proposed principle-aware system leverages Large Language Model (LLM)-based MAS to conduct scientific discovery through strategic hypothesizing. The framework is comprised of two core, interconnected components: (1) an MAS that executes a Hypothesis-Validation loop, and (2) PiFlow, which serves as the strategic director for this loop.
Definition 3.1 (Scientific Principles).
The scientific principles are foundational concepts, established laws, or patterns that explain phenomena within a specific scientific domain. These principles serve as high-level conceptual building blocks from which specific, testable hypotheses can be derived. This conceptualization aligns with broader discussions on the nature of scientific knowledge Poincar1906ScienceAH .
Hypothesis-Validation Loop. As depicted in the right-hand side of Figure 2, the Hypothesis-Validation loop constitutes the core operational cycle and incorporates two LLM-based agents: Hypothesis Agent ($\mathcal{A}_{H}$) and Experiment Agent ($\mathcal{A}_{E}$), detailed below:
  1. (a)
    Hypothesis. Initially, a set of candidate principles $\mathcal{P}=\{p_{1},p_{2},p_{3}\}$ (see Definition 3.1)potentially proposed by domain experts or extracted by LLMs, is established. In each iteration $t$, $\mathcal{A}_{H}$ proposes a testable hypothesis, $h_{t}$, grounded in a selected principle $p_{i}\in\mathcal{P}$. The proposal is supported by structured rationales (comprising major and minor premises, following (Eisape2023ASC, )), rather than unconstrained association.
  2. (b)
    Validation. Subsequently, $\mathcal{A}_{E}$ rigorously validates $h_{t}$ using an experimental tool, denoted $f^{*}(\cdot)$, which yields a quantitative outcome $y_{t}=f^{*}(h_{t})$ (e.g., property value of a material). This outcome serves as feedback for subsequent hypothesizing.
This iterative process progressively establishes a record of principle-outcome pairs, $\mathcal{T}_{t}=\{\left<p_{k},y_{k}\right>\}_{k=1}^{t}$, linking each hypothesized principle $p_{k}$ to its observed experimental outcome $y_{k}$. While this loop systematically generates valuable evidence in trajectory $\mathcal{T}_{t}$, the selection of potential principles for subsequent hypotheses can lack strategic direction if not externally guided. This may lead to inefficient exploration of the principle space or premature convergence on suboptimal findings.
Hypothesis steering with PiFlow. To address the potential inefficiencies in principle selection and to instill strategic direction, the PiFlow component is introduced. It leverages the dynamically growing set of principle-outcome pairs, $\mathcal{T}_{t}$ as its primary input. Two steps are conducted to steer the hypothesizing:
  1. (a)
    High-potential principle acquisition. After an initial phase of evidence collection, during which $\mathcal{T}_{t}$ is populated, PiFlow activates its core mechanism: an adversarial Min-Max optimization (detailed in Section 3.3). This optimization process analyzes $\mathcal{T}_{t}$ to identify a principle, $p^{*}$ (e.g., $p_{2}$ in Figure 2), predicted to have the highest potential for advancing the scientific inquiry.
  2. (b)
    Principle steering for hypothesizing. Following the identification of the highest-potential principle $p^{*}$ by the Min-Max optimization, PiFlow assigns a potential score to all principles in $\mathcal{P}$. This scoring enables a threshold-based partitioning detailed in Appendix C.1: principles with high scores (like $p^{*}$, e.g., $p_{2}$) drive refinement; those with medium scores trigger further validation ($p_{2}\rightarrow p_{2}^{\prime}$); and low scores prompt exploration of new conceptual areas.
This closed-loop feedback mechanism ensures that the Hypothesis-Validation cycle is continuously and adaptively steered by strategic insights derived from the system’s cumulative experience, as embodied in $\mathcal{T}_{t}$.
Plug-and-Play Modularity of PiFlow. The hypothesis steering mechanism above, which includes PiFlow and its interfacing Planner agent $\mathcal{A}_{P}$, has been intentionally engineered as a modular, Plug-and-Play system. This architectural choice ensures that principle-aware guidance can be readily integrated to enhance MAS engaged in scientific discovery as a part of prompts. The Planner agent, $\mathcal{A}_{P}$, serves as the critical adaptable interface, channeling analytically derived principle suggestions from PiFlow directly into the operational workflow of a host MAS. This adaptability positions PiFlow as a pivotal enhancement for automated scientific inquiry.
Figure 2: Overview of the PiFlow Architecture for Scientific Discovery. The PiFlow component utilizes Min-Max optimization to strategically select and direct high-potential principles to the Planner agent. The Planner, in turn, guides the Hypothesis-Validation loop, where agents iteratively generate hypotheses $h_{t}$ from principles $p_{t}$ at step $t$, execute experiments, and refine understanding. This iterative process is designed to efficiently navigate the discovery landscape.
3.3 Min-Max Optimization in PiFlow
In our proposed method, strategic principle selection is achieved through a Min-Max optimization, as presented in Eq. 1. This approach is designed to balance the exploitation of established, high-potential principles (which then guide the formulation of specific hypotheses) with the exploration of novel ones, while explicitly incorporating information acquisition efficiency.
$\min_{\pi\in\Pi}\max_{f^{*}\in\mathcal{F}}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))-\lambda\cdot I(h_{t};f^{*}|H_{t-1})\right]$ (1)
where $\pi\in\Pi$ represents the decision-making policy for selecting principles and $f^{*}\in\mathcal{F}$ is the evaluation function (e.g., an experimental tool) that characterizes the quantitative outcome yielded from hypothesis $h_{t}\in\mathcal{H}$, where $\mathcal{H}$ is the hypothesis space. Over $T$ iterations, the policy $\pi$ aims to minimize the objective function in Eq. 1. This objective strategically balances: (1) the summation for cumulative regret, encouraging exploitation of known high-potential principles, and (2) the mutual information, thereby effectively maximizing information gain to foster exploration. This structure allows PiFlow to navigate the complex trade-offs between these two goals.
Minimizing Cumulative Regret (Exploitation). The first term, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, represents the cumulative regret over $T$ iterations. Here, $v^{*}$ is a theoretical optimal outcome value, and $f^{*}(h_{t})$ is the outcome from hypothesis $h_{t}$. By minimizing this term, the policy $\pi$ is driven to exploit known high-potential principles and hypotheses to achieve outcomes as close as possible to the optimum $v^{*}$. This encourages the refinement and validation of promising avenues.
Maximizing Information Gain (Exploration). The second term $-\lambda\cdot I(h_{t};f^{*}|H_{t-1})$ promotes exploration. The policy $\pi$ seeks to minimize this term, which is equivalent to maximizing the mutual information $I(h_{t};f^{*}|H_{t-1})$. This mutual information quantifies the expected reduction in uncertainty about the true evaluation function $f^{*}$ upon observing the outcome of hypothesis $h_{t}$, given all past observations $H_{t-1}=\{\left(h_{m},y_{m}\right)\}_{m=1}^{t-1}$. The trade-off parameter $\lambda>0$ controls the balance between minimizing regret (exploitation) and maximizing information gain (exploration). A larger $\lambda$ places more emphasis on information acquisition.
Remark (The dependencies of $f^{*}$ in $I(h_{t};f^{*}|H_{t-1})$) . The informativeness of a proposed hypothesis $h_{t}$ is inherently dependent on the nature of the true underlying evaluation function $f^{*}$. The adversarial nature of the $max_{f^{*}}$ operator means that the policy $\pi$ must select hypotheses $h_{t}$ that are expected to be informative even if $f^{*}$ were to manifest in a way that makes $h_{t}$ minimally revealing, thus ensuring robustness in information acquisition.
In summary, the Min-Max adversarial formulation underpinning PiFlow provides strong theoretical guarantees, notably sublinear regret bounds (formalized in Theorem 1, with proof in Appendix B). Importantly, PiFlow’s operational behavior, which involves practical approximations of this Min-Max solution, demonstrates consistent alignment with these theoretical expectations, as empirically validated in Section 5.3.
Theorem 1 (Informal).
The Min-Max optimization in PiFlow formulates a trade-off between exploitation (minimizing regret) and exploration (maximizing information gain). Under conditions of finite entropy $H(f^{*})$ and bounded evaluation function $f^{*}$, this optimization provides two key theoretical guarantees: (1) As information gain decreases, the expected regret also decreases; (2) the cumulative regret grows at a sublinear rate of $\mathcal{O}\left(\sqrt{T}\right)$.
4 Experiment
4.1 Settings
Simulating hypothesis validation is crucial for our experiments. However, designing individual simulation environments for each experiment is costly. Therefore, we employ surrogate models deployed locally, behaving like a HTTP-based tool for the experiment agent. In our experiments, the task can be “Find a molecule candidate with the highest bio-activity.” To address such tasks, the hypothesis agent proposes hypotheses and the experiment agent validate them using the tool, aiming to identify optimal candidates.
To ensure a fair comparison and focus on core capabilities, all agents utilize QwenMax (Yang2024Qwen2TR, ) as the base LLM and are prohibited from accessing external search tools. The implementation details are in Appendix C.1, and role-playing prompts for all agents are provided in Appendix E. Each experiment was repeated three times with different random seeds, and we report the mean and standard deviation (std) of the relevant metrics for each method.
4.2 Experimental Scenarios
Nanohelix Optimization (NHO, see Appendix D.1). We developed a high-fidelity surrogate model ($r^{2}=0.98$) to predict g-factor values of a nanohelix material. The model was trained on DFT-simulated data Wu2025MachineLS , enabling efficient exploration of the four-dimensional parameter space (radius of fiber and helix, number of turns, etc.), describing the physical structure of a standard helix.
Molecular Bio-activity Optimization (MBO, see Appendix D.2). We constructed a surrogate model ($r^{2}=0.91$) to predict molecular bio-activity using 50,000 molecule-activity pairs sampled from ChEMBL35 Zdrazil2023TheCD . This model maps SMILES representations directly to pChEMBL values, facilitating rapid in silico screening of candidate therapeutics.
Superconductor Optimization (SPO, see Appendix D.3). We constructed a surrogate model ($r^{2}=0.91$) to predict a superconductor’s $T_{c}$ using 26,321 records of superconductor data publicly available (Hamidieh2018ADS, ). The model learns the map between the compositional features of superconductor to the $T_{c}$ value, facilitating the discovery of room-temperature superconducting materials.
4.3 Baselines
To evaluate PiFlow’s specific guidance strategy under uncertainty, we therefore benchmark against the following baselines: (1) Reasoning and Acting (ReAct) (Yao2022ReActSR, ). ReAct enables agents to iteratively formulate hypotheses, design/execute experiments, and interpret results, representing a foundational approach to the scientific method loop. (2) Meta Plan Optimization (MPO) (Xiong2025MPOBL, ). MPO utilizes an LLM planner that generates guidance based on past exploration experiences. This offers a contrast to PiFlow’s distinct, explicit mechanism for principled uncertainty reduction. (3) Vanilla Agent System (Vanilla). A Vanilla agent, where a single LLM conducts the discovery process independently, serves as a fundamental benchmark to illustrate the challenges of unguided exploration. While many other advanced agent frameworks exist, their core focus often differs. For instance, Agent-Oriented Planning (AOP) (Li2024AgentOrientedPI, ) is optimized for tasks with known structures rather than initial epistemic uncertainty, and Reason for Future, Act for Now (RAFA) (Liu2023ReasonFF, ) prioritizes predefined task completion over the iterative evidence building from scratch that PiFlow targets. Our evaluation against these baselines is designed to highlight PiFlow’s advantages in discovery efficiency.
4.4 Evaluation Metrics
We employ two metrics to comprehensively evaluate the performance of PiFlow, as detailed below.
Solution Quality (SQ). We measure the optimal objective value with a percentage relative to the theoretical maximum value $\mu_{\text{absolute}}$, denoted as:
$\mathrm{SQ}=\frac{\max\{y_{k}|\left<p_{k},y_{k}\right>\in\mathcal{T}_{t}\}}{\mu_{\text{absolute}}}\times 100\%.$ (2)
Specifically, for nanomaterial chirality with g-factor value, the theoretical maximum is reported as $\mu_{\text{absolute}}^{\text{g-factor}}=2.0$, indicating a total selectivity for CP light (Greenfield2021PathwaysTI, ), the larger the g-factor, the stronger the chirality; for molecular bio-activity with pChEMBL value, a strict threshold of $\mu_{\text{absolute}}^{\text{pchembl}}=6.5$ has been reported (Lenselink2017BeyondTH, ) for lead compound optimization stage, larger value indicates a better bio-activity and vice versa. for superconductivity Tc, the reference value is set to be $25^{\circ}\mathrm{C}$ , which is equal to 298.15K, denoted as $\mu_{\text{absolute}}^{\text{Tc}}=298.15K$.
Area Under the Curve (AUC). Exploration efficiency requires both (1) rapid convergence and (2) high objective values. We quantify these two factors by defining the AUC metric. Given the trajectory $p_{k},y_{k}\right>\in\mathcal{T}_{t}$ across $t$ steps, AUC is computed via the trapezoidal rule. For meaningful comparisons, we normalize by the maximum possible area:
$\mathrm{AUC}=\frac{\sum_{i=1}^{t-1}\frac{y_{i}+y_{i+1}}{2}}{\mu_{\text{absolute}}\cdot(t-1)}\times 100\%$ (3)
This normalized metric captures the ability to maintain high performance throughout the optimization trajectory, with higher values indicating more efficient exploration.
5 Results
5.1 Performance Comparison
We use SQ to compare the overall capability of reaching the objective solution, and AUC to assess the efficiency, reflecting progress towards better outcomes over the exploration process.
Table 1 demonstrates that PiFlow achieves significant improvements over all baselines across three benchmarks of NHO, MBO, and SPO. In terms of achieving the target property, measured by SQ, PiFlow consistently leads, outperforming ReAct, MPO, and Vanilla systems by an average of approximately 207.6%, 34.1%, and 94.1%. Furthermore, PiFlow demonstrates superior exploration efficiency, evidenced by its leading AUC scores, translating to average AUC improvements of approximately 146.5% over ReAct, 73.6% over Vanilla and 54.0% over MPO. Additionally, we compare the exploration trajectory regarding the experiment outcome, as presented in Figure 3. It shows that, PiFlow is always the fastest one in reaching the best solution while exploring. Although there are some fluctuations in the process, the general upward trend continues to surpass other methods.
These significant advancements in both SQ and AUC strongly support our central claim: the strategic and principled guidance provided by the PiFlow component, rooted in its principle-aware optimization, effectively steers the multi-agent hypothesis-validation loop towards more rapid and fruitful scientific discovery.
Table 1: Comparisons between PiFlow and baselines.
Method NHO (g-factor) MBO (pChEMBL) SPO ($T_{c}$)
AUC (%) SQ (%) AUC (%) SQ (%) AUC(%) SQ (%)
ReAct 35.85 $\pm$7.12 41.96 $\pm$0.82 29.61 $\pm$9.74 43.11 $\pm$9.98 5.29 $\pm$0.69 6.41 $\pm$0.96
Vanilla 35.96 $\pm$22.38 46.76 $\pm$7.29 29.71 $\pm$12.66 49.22 $\pm$8.30 11.39 $\pm$11.33 14.16 $\pm$13.37
MPO 43.99 $\pm$2.79 51.29 $\pm$7.77 31.18 $\pm$9.12 57.28 $\pm$5.85 12.68 $\pm$7.75 33.20 $\pm$23.75
PiFlow (ours) 63.51 $\pm$11.18 76.82 $\pm$4.54 46.11 $\pm$16.25 84.55 $\pm$29.63 21.51 $\pm$2.80 34.85 $\pm$1.19
Figure 3: Trajectory comparisons for different optimization methods.
5.2 Ablation Study
We conduct several ablations to evaluate the components of PiFlow. The studies investigate the impact of PiFlow itself, the choice of foundation LLMs, the influence of a key hyperparameter $\lambda$, which is for balancing exploitation and exploration, and the effect of the Thought mode in LLM reasoning. For these studies, conducted on the NHO task, performance is evaluated based on the same metrics, AUC (%) and SQ (%).
Impact of PiFlow Integration (Plug-and-Play). To isolate the direct benefit of PiFlow, we compared the performance of two different LLMs, GPT4.1-mini and Qwen3-32B, w/ and w/o the PiFlow component. This is achieved by only including or excluding the steered principle with one of the three specified action types to the Planner Agent via prompt, which then directs subsequent hypothesizing and validation.
Table 2: Ablation Study with/without PiFlow
Method/Setting AUC (%) SQ (%)
GPT4.1-mini w/ PiFlow 41.68 $\pm$17.91 66.38 $\pm$14.90
w/o PiFlow 37.12 $\pm$16.04 40.14 $\pm$13.97
Qwen3-32B w/ PiFlow 37.51 $\pm$7.70 58.76 $\pm$6.18
w/o PiFlow 27.04 $\pm$21.50 54.84 $\pm$24.41
As shown in Table 2, for GPT4.1-mini, integrating PiFlow increases the AUC from 37.12% to 41.68% and substantially boosted the SQ from 40.14% to 66.38%. This represents an approximate 12.3% improvement in AUC and a significant 65.4% improvement in SQ. Similarly, for the Qwen3-32B model, the inclusion of PiFlow improves AUC from 27.04% to 37.51% (a 38.7% increase) and SQ from 54.84% to 58.76% (a 7.1% increase). $\diamond$ Takeaway: These findings underscore the significant positive impact of PiFlow on both AUC and SQ, confirming its effectiveness as a Plug-and-Play enhancement for MAS in scientific discovery.
Influence of Foundation Models. The choice of underlying LLMs may significantly affect the reasoning process, thereby affecting the performance. To evaluate the model influence while balancing the agent functions, we replace LLMs of $\mathcal{A}_{H}$ and $\mathcal{A}_{P}$ evaluation. The $\mathcal{A}_{E}$responsible for tool usage, consistently used QwenMax to ensure functional tool interaction. We evaluate several state-of-the-art LLMs (e.g., Claude-3.7-sonnet anthropic2025 , GPT4.1-mini openai2025 , Gemini-2.5-pro-exp-0325 google2025 , Qwen3-32B and QwenMax yang2024qwen2 ) on the NHO task, and the results are summarized in Table 3.
Table 3: Model Performance Comparison
Method/Setting AUC (%) SQ (%)
Claude-3.7-sonnet 38.60 $\pm$4.30 78.50 $\pm$3.74
GPT4.1-mini 41.68 $\pm$17.91 66.38 $\pm$14.90
Gemini-2.5-pro-exp-03-25 28.43 $\pm$13.81 69.64 $\pm$17.10
Qwen3-32B 37.51 $\pm$7.70 58.76 $\pm$6.18
QwenMax 63.51 $\pm$11.18 76.82 $\pm$4.54
Among these models, QwenMax demonstrates the highest AUC at 63.51% and a strong SQ of 76.82%. Claude-3.7-sonnet achieves the highest SQ at 78.50% with an AUC of 38.60%. Other models like GPT4.1-mini (AUC 41.68%, SQ 66.38%) and Qwen3-32B (AUC 37.51%, SQ 58.76%) also show competent performance. Gemini-2.5-pro-exp-03-25 yields an AUC of 28.43% and an SQ of 69.64%. $\diamond$ Takeaway: The performance variation across different models highlights the importance of model selection for specific tasks within the PiFlow-enhanced MAS.
Sensitivity to Hyperparameter $\lambda$. We investigate the impact of $\lambda$ in Eq. 1. The specific role of $\lambda$ is to balance exploration and exploitation, with larger $lambda$ value places greater emphasis on exploration. The results conducted on the NHO task using QwenMax model for different values of $\lambda$ are detailed in Table 4.
Table 4: Lambda Ablations
Setting AUC (%) SQ (%)
$\lambda=0.1$ 41.32 $\pm$5.90 66.45 $\pm$9.49
$\lambda=0.3$ 44.28 $\pm$2.83 66.43 $\pm$9.28
$\lambda=0.5$ 32.99 $\pm$11.16 59.02 $\pm$3.56
$\lambda=0.7$ 40.50 $\pm$2.89 56.40 $\pm$4.79
$\lambda=0.9$ 34.57 $\pm$10.33 62.49 $\pm$8.38
From the table, AUC varied with different $\lambda$ values, peaking at 44.28% when $\lambda=0.3$. The SQ remained relatively high for $\lambda=0.1$ (66.45%) and $\lambda=0.3$ (66.43%), but showed more variability with other settings. For instance, $\lambda=0.5$ resulted in a lower AUC (32.99%) and SQ (59.02%). While increasing $\lambda$, AUC tends to decrease, as a stronger emphasis on exploration can lead to more varied hypothesis selection. These results suggest that the system’s performance, particularly its exploration efficiency, is sensitive to the choice of $\lambda$, with $\lambda=0.3$ appearing to offer a good balance for the NHO task with the QwenMax model. $\diamond$ Takeaway: The results suggest that $\lambda$ indeed effects the performance in both AUC and SQ, fine-tuning $\lambda$ can be crucial for optimizing performance.
Effect of Thought Mode. We also conducted an ablation study on an internal thought mode of the LLM (referred to as Think in Table 5) on the NHO benchmark. This is for agents based on Qwen3-32B and Qwen3-8B models, which support ON/OFF <think> token generation with system prompt including or excluding /no_think. This mode is intended to enable more explicit reasoning steps.
Table 5: Effect of Thinking.
Method/Setting AUC (%) SQ (%)
Qwen3-32B w/ Think 37.51 $\pm$7.70 58.76 $\pm$6.18
w/o Think 45.51 $\pm$11.19 68.86 $\pm$17.67
Qwen3-8B w/ Think 30.55 $\pm$27.45 54.49 $\pm$22.25
w/o Think 42.09 $\pm$16.55 61.59 $\pm$16.43
Interestingly, for both Qwen3-32B and Qwen3-8B, disabling the Thought Mode led to improved performance. For Qwen3-32B, removing the Thought Mode increased AUC from 37.51% to 45.51% and SQ from 58.76% to 68.86%. Similarly, for Qwen3-8B, performance without the Thought Mode was better, with AUC rising from 30.55% to 42.09% and SQ increasing from 54.49% to 61.59%. $\diamond$ Takeaway: These counter-intuitive results suggest that the current implementation or utility of this specific Thought Mode may introduce overhead or misguide the agents in these contexts, warranting further investigation into its design and application.
5.3 Theoretical Alignment
To empirically validate PiFlow’s theoretical guarantees (Theorem 1), we analyze key aspects of its exploration dynamics, including (1) the bound of average regret in PiFlow and (2) the relationship between regret and information gain, as shown in Figure 4.
(Theoretical Prediction 1) Average regret decay with $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$. Figure 4a presents the average regret as a function of iterations on a log-log scale. The empirical results demonstrate a strong concordance with the theoretical prediction of $\mathcal{O}(\sqrt{T})$ sublinear cumulative regret. This is evidenced by average regret trajectories adhering to the $T^{-0.5}$ decay (fitted $c\cdot T^{-0.5}$ with $c=1.37,R^{2}=0.96$), a pattern most runs consistently follow. Notably, one run (ID=1), after an initial sharp regret decrease, shows a transient increase before resuming decay. This illustrates PiFlow’s robust exploration avoiding potential local optima with a characteristic of its Min-Max strategy.
(Theoretical Prediction 2) As information gain decreases, the expected regret also decreases. Theorem 1 also posits that decreasing information gain correlates with decreasing expected regret. Figure 4b empirically verifies this with PiFlow. The scatter plot of regret versus information gain (points colored by iteration) reveals a clear positive association, confirmed by a fitted trend line: lower information gain generally corresponds to lower regret. Early iterations (higher information gain and regret) transition to later iterations (lower information gain and regret). This empirical trend strongly supports the theoretical analysis, showcasing PiFlow’s effective shift from exploration to exploitation.
(a) Average Regret Dynamics
(b) Regret vs. Information Gain
(c) NHO Exploration Trajectory
Figure 4: Empirical Validation of PiFlow’s Theoretical Alignment. (a) Average regret over iterations aligns with the theoretical sublinear decay; (b) Scatter plot with positive trend of observed regret versus information gain, supporting theoretical expectations; (c) An exemplary PiFlow exploration trajectory.
Empirical Validation of Exploration Strategy on NHO Landscape. To further illustrate PiFlow’s operational dynamics, Figure 4c visualizes an exemplary exploration trajectory on the NHO task. The objective is to identify nanohelix parameters yielding the maximum property value. The background displays the NHO objective landscape by Principal Component Analysis (PCA), where contours indicate g-factor values fitted over the source data.
Aligning with its robust design, PiFlow’s trajectory shows initial broad exploration (iterations 1-16), then successful navigation of a low-quality “valley” (iterations 16-21) crucial for avoiding local optima, before efficiently converging to a high-g-factor region and identifying near-optimal parameters (iterations 21-24).
$\diamond$ Takeaway: The above regret analysis and visualized behavior empirically support the claim that PiFlow’s principle-aware mechanism translates its theoretical guarantees into effective and efficient scientific discovery.
6 Limitation
While PiFlow shows notable improvements through its principled Min-Max optimization, its practical implementation approximates a key theoretical component. This means the current system may not fully capture all nuances of true, model-based information gain, especially the direct adversarial interplay with all possible manifestations of the unknown evaluation function $f^{*}$ from the theoretical objective. Future research could explore more direct estimations of mutual information for this heuristic within the PiFlow framework to potentially further enhance its strategic guidance.
7 Conclusion
In conclusion, we propose PiFlow, a Plug-and-Play module to strategically guide the Hypothesis-Validation loop through a principle-steering mechanism to address challenges of aimless hypothesis searching and unclear connections between hypotheses and evidence. Our approach utilizes a Min-Max optimization that explicitly balances exploitation of high-potential principles with exploration of novel hypotheses, guaranteed by a sublinear average regret bound. Extensive experiments demonstrate that PiFlow can adaptively navigate complex hypothesis spaces without premature convergence on suboptimal solutions, yielding significant improvements over baseline methods. In addition, PiFlow enhanced process demonstrates remarkable ability in avoiding local optima, yet keeping structured rationales for better interpretability. We hope PiFlow can contribute to advancing automated scientific discovery, inspiring further exploration and innovation.
Appendix A Comparison of Limitations
While existing methods of LLM-based agent systems demonstrate significant impact in advancing scientific discovery, they limited to direct application in other domains without any changes of architectures and prompts.
Table 6: Comparison of Generalization Limitations in LLM-Agent Systems
Method Name Core Implementation Adaptability
The AI Scientist [30] Generating novel research ideas, writes code, executes experiments, visualizes results, describes its findings by writing a full scientific paper AI research
Agent Laboratory [45] Accepting a human-provided research idea and progressing through three stages – literature review, experimentation, and report writing to produce comprehensive research outputs, including a code repository and a research report AI research
CellAgent [52] Constructing LLM-driven biological expert roles - planner, executor, and evaluator - each with specific responsibilities scRNA-seq Analysis
DrugAgent [28] Employing an LLM Planner that formulates high-level ideas and an LLM Instructor that identifies and integrates domain knowledge when implementing those ideas Drug discovery
IAN [35] Leveraging popular pathway and regulatory datasets for protein-protein interactions to perform analysis through a LLM-based multi-agent architecture Protein-protein interactions analysis
Alloy Design AI [14] Automating alloy design and discovery using physics-aware multimodal multi-agent AI Alloy design and discovery
Inorganic Materials AI [47] Accelerating inorganic materials design using generative AI agents. Inorganic materials design
LIDDIA [5] Language-based intelligent agent for drug discovery tasks. Drug discovery
dZiner [3] Rational inverse design of materials facilitated by AI agents. Inverse design of materials
MOOSE-Chem [58] Utilizing Large Language Models for rediscovering unseen chemistry scientific hypotheses. Chemistry hypothesis discovery
OmniScience [39] Domain adaptive pretraining, instruction tuning and reasoning-based knowledge distillation Scientific reasoning
Organic Semiconductor AI [65] Large Language Model-Based AI Agent for research in organic semiconductor devices Organic semiconductor device research
DrugAgent (Explainable) [21] Explainable drug repurposing agent with Large Language Model-based reasoning Drug repurposing
ProtAgents [11] Protein discovery via large language model multi-agent collaborations combining physics and machine learning Protein discovery
To illustrate the challenge of limited generalization, Table 6 outlines several representative LLM-based agent methods. It highlights how their core implementations and prompt engineering strategies are often tightly coupled to specific scenarios, thereby restricting their out-of-the-box applicability to new or diverse problem domains. This tight coupling necessitates significant modifications when attempting to transfer these systems to different contexts, underscoring a key area for future research and development in creating more adaptable and universally applicable AI agents.
Appendix B The Proof of the Convergence
We recall the elements here and formally proof the convergence along with boundary of the system. Here we denote $\pi$ is the language model policy from policy space $\Pi$, $f^{*}$ is the acquisition function from function space $\mathcal{F}$, $h_{t}$ is the hypothesis at time step $t$, $H_{t-1}=\{h_{1},h_{2},\dots,h_{t-1}\}$ is the history of hypotheses, $v^{*}$ is the SQ achievable by any hypothesis, and $I(h_{t};f^{*}\mid H_{t-1})$ is the conditional mutual information.
According to the original formulation (Eq. 1), the cumulative regret can be expressed as
$R_{T}(\pi,f^{*})=\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))\right]$
and the cumulative information gain is
$IG_{T}(\pi,f^{*})=\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}I(h_{t};f^{*}|H_{t-1})\right]$
Information gain approaches zero. With information theory, the mutual information can be written as:
$I(h_{t};f^{*}|H_{t-1})=H(f^{*}|H_{t-1})-H(f^{*}|H_{t-1},h_{t})$
A critical property for convergence is that the total information gain is bounded,
$\sum_{t=1}^{\infty}I(h_{t};f^{*}\mid H_{t-1})\leq H(f^{*})<\infty$
This follows from the chain rule of mutual information,
$\sum_{t=1}^{T}I(h_{t};f^{*}\mid H_{t-1})=I(H_{T};f^{*})\leq H(f^{*})$
Since the entropy $H(f^{*})$ is finite, the cumulative information gain is bounded regardless of how many steps $T$ we take. This implies that:
$\lim_{t\to\infty}I(h_{t};f^{*}\mid H_{t-1})=0.$
In other words, the marginal information gained from each new hypothesis must eventually approach zero.
As information gain decreases, the expected regret also decreases. As we mentioned before, the regret $R_{T}(\pi,f^{*})$ is defined as:
$R_{T}(\pi,f^{*})=\mathbb{E}_{T}[v^{*}-f^{*}(h_{t})].$
Now apply Jensen’s Inequality, let $X=v^{*}-f^{*}(h_{t})$, we have
$\mathbb{E}[X^{2}\mid H_{t-1}]=\mathbb{E}[(v^{*}-f^{*}(h_{t}))^{2}\mid H_{t-1}].$
Use $\phi(x)=x^{2}$, as it is convex, giving
$(\mathbb{E}[X\mid H_{t-1}])^{2}\leq\mathbb{E}[X^{2}\mid H_{t-1}],$
take square roots:
$\mathbb{E}[v^{*}-f^{*}(h_{t})\mid H_{t-1}]\leq\sqrt{\mathbb{E}[(v^{*}-f^{*}(h_{t}))^{2}\mid H_{t-1}]}.$
Now we deal with the second moment of the regret. As the $v^{*}$ is constant, therefore, with the variance formula, we have
$\mathbb{E}\left[(v^{*}-f^{*}(h_{t}))^{2}\mid H_{t-1}\right]=\text{Var}(f^{*}(h_{t})\mid H_{t-1})+\left(v^{*}-\mathbb{E}[f^{*}(h_{t})\mid H_{t-1}]\right)^{2}$
Both terms (variance and bias) are non-negative, and our goal is to bound this expression. The variance term $\text{Var}(f^{*}(h_{t})\mid H_{t-1})$ captures the uncertainty in $f^{*}(h_{t})$ given the history.
Since the information theory proofed that, the variance of a function can be bounded by mutual information, akin to entropy bounds $H(f^{*})\leq\log(|\mathcal{F}|)$, for the first term, we have
$\text{Var}(f^{*}(h_{t})\mid H_{t-1})\leq c\cdot I(h_{t};f^{*}\mid H_{t-1}),$
where $c$ is constant that depends on the range of $f^{*}$, this follows because the mutual information bounds the expected variance of conditional expectations.
Since we can direct view the $f^{*}(h_{t})$ as a random variable over the joint distribution of $f^{*}$ and $h_{t}$, we can apply a concentration inequality. A standard result in information-directed sampling states that for a bounded random variable like $f^{*}(h_{t})$ here, the second moment of the regret $(v^{*}-f^{*}(h_{t}))$ can be bounded as:
$\mathbb{E}\left[(v^{*}-f^{*}(h_{t}))^{2}\mid H_{t-1}\right]\leq c\cdot I(h_{t};f^{*}\mid H_{t-1}),$
where the constant $c$ depends on $|\mathcal{F}|$.
Finally, we get the inequality of $R_{T}(\pi,f^{*})$:
$R_{T}(\pi,f^{*})\leq\sqrt{\mathbb{E}_{T}\left[(v^{*}-f^{*}(h_{t}))^{2}\mid H_{t-1}\right]}\leq c\cdot\sqrt{I(h_{t};f^{*}\mid H_{t-1})}$
This inequality demonstrates that as the information gain $I(h_{t};f^{*}\mid H_{t-1})$ decreases, the expected regret also diminishes.
The cumulative regret grows at a rate $O(\sqrt{T})$. Based on the inequality in the last step and to find the cumulative regret bound, we need to sum this inequality over all time steps:
$\sum_{t=1}^{T}R_{t}(\pi,f^{*})\leq c\cdot\sum_{t=1}^{T}\sqrt{I(h_{t};f^{*}|H_{t-1})}$
Applying the Cauchy-Schwartz inequality:
$\sum_{t=1}^{T}\sqrt{I(h_{t};f^{*}|H_{t-1})}\leq\sqrt{T\cdot\sum_{t=1}^{T}I(h_{t};f^{*}|H_{t-1})}$
Since we’ve already established that the total information gain is bounded:
$\sum_{t=1}^{T}I(h_{t};f^{*}|H_{t-1})\leq H(f^{*})$
We can substitute this bound:
$\sum_{t=1}^{T}\sqrt{I(h_{t};f^{*}|H_{t-1})}\leq\sqrt{T\cdot H(f^{*})}$
Therefore, the cumulative regret is bounded by:
$\sum_{t=1}^{T}R_{t}(\pi,f^{*})\leq c\cdot\sqrt{T\cdot H(f^{*})}$
This demonstrates that the cumulative regret grows at a rate of $\mathcal{O}(\sqrt{T})$, which is sublinear in $T$. This result implies that while the total regret increases with step, the average regret per time step decreases at a rate of $\mathcal{O}(\frac{1}{\sqrt{T}})$.
Appendix C Implementation Details
C.1 Algorithm of PiFlow
1
2
Input: $\mathcal{T}_{t}$, and $\lambda_{factor}$
Output: $suggestion$ (string action recommendation)
3
4if $|\mathcal{F}|<3$ then
       return “Initialize one principle to explore. ”
        // Not enough principles
5
6 end if
7
[…]
  // $y_{t}$ in $\mathcal{T}_{t}$ as exploitation score for each $p_{t}$
9 $S_{exploitation}\leftarrow{compute\_exploitation\_scores}(\mathcal{T}_{t})$
  // Embedding differences to others with cosine similarity
10
11for $i\leftarrow 1$ to $|\mathcal{T}|$ do
12       $S_{final}[i]\leftarrow(1-\lambda)\cdot S_{exploration}[i]+\lambda\cdot S_{exploitation}[i]$
13
14 end for
15
[…]
17 $p_{best}\leftarrow\mathcal{T}[i_{best}]$
18 $best\_exploitation\_score\leftarrow S_{exploitation}[i_{best}]$
  // $\triangleright$ Score for the chosen principle
19
20if $best\_exploitation\_score>0.7$ then
[…]
22       $suggestion\leftarrow\text{``Focus on refining: }\frown p_{best}.content\frown\text{High potential.''}$.”
        // $\triangleright$ REFINE
23
24 else if $best\_exploitation\_score>0.4$ then
[…]
26       $suggestion\leftarrow\text{``Validate: }\frown p_{best}.content\frown\text{Moderate potential.''}$.”
        // $\triangleright$ VALIDATE
27
28 else
[…]
30       $suggestion\leftarrow\text{``Explore alternatives: }\frown p_{best}.content\frown\text{Low current potential''}$”
        // $\triangleright$ EXPLORE
31
32 end if
33
34return $suggestion$
Algorithm 1 Algorithm of the PiFlow Policy
Appendix D Experiment Environment
D.1 Nanohelix Optimization
D.1.1 Problem Statement
Nanohelices are helical nanostructures with unique physical properties that make them valuable for applications in electronics, photonics, and magnetism. Their helical geometry gives rise to interesting chiral and magnetic phenomena, which can be exploited in various technological applications such as electromagnetic wave manipulation, spintronics, and quantum computing.
Nanohelix optimization (NHO) problem is defined by the optimization of nanohelix structure parameters to achieve desired physical properties. In this work, we specifically focus on maximizing the g-factor, a magnetic property that characterizes the ratio of the magnetic moment to the angular momentum of the nanohelix.
The nanohelix structure is characterized by four key geometric parameters:
Fiber-radius ($r_{f}$, nm): Radius of the actual fiber/wire that forms the helix structure. The values for this parameter range from 20 nm to 60 nm, with 10 evenly spaced values.
Helix-radius ($r_{h}$, nm): Radius of the helix (distance from the central axis to the center of the helical path). The values for this parameter range from 20 nm to 90 nm, with 10 evenly spaced values.
Number of turns ($n_{t}$, dimensionless): Number of complete turns in the helix. The values for this parameter range from 3 to 10, with 10 evenly spaced values.
Pitch ($p$, nm): Axial distance between adjacent turns. The values for this parameter range from 60 nm to 200 nm, with 10 evenly spaced values.
Mathematically, we can formulate the optimization problem as:
$\arg\max_{\theta\in\Theta}\quad f(\theta)$
Figure 5: The $r^{2}$ of the surrogate model.
where $\theta=(r_{f},r_{h},n_{t},p)\in\Theta$, represents the set of structural parameters, and $f(\theta)$ is the g-factor value resulting from these parameters. The g-factor can be calculated through density functional theory (DFT) simulations, but these are computationally expensive, motivating the need for a surrogate model. The modification of these parameters, as an example, can be seen at Figure 6.
Figure 6: The demonstrated changes of helix parameters and their property value (g-factor).
D.1.2 Development
The dataset contains 6300 records of nanohelix structural parameters with corresponding g-factor values. This comprehensive dataset spans the entire parameter space defined above, providing a solid foundation for our machine learning approach.
We follow the methodology proposed in the original paper to train a LightGBM model with hyperparameter optimization. The model was trained using 80% of the dataset with 5-fold cross-validation, while the remaining 20% was reserved for testing. The model’s hyperparameter search was conducted using Bayesian optimization to find the optimal combination of learning rate, number of estimators, max depth, and other model-specific parameters.
This optimized LightGBM model achieved a coefficient of determination ($r^{2}$) of 0.9802, as shown in Figure 5, indicating that the model explains 98.02% of the variance in the g-factor prediction given any structural parameters. This high level of accuracy enables reliable exploration of the parameter space without requiring computationally expensive DFT simulations for each parameter combination.
D.2 Bio-activity Optimization
D.2.1 Problem Statement
Molecular bio-activity refers to the ability of a chemical compound to interact with biological targets, such as proteins, enzymes, or receptors, and induce a biological response. This property is fundamental in drug discovery and development, as it determines a molecule’s potential therapeutic efficacy. The strength of this interaction is often quantified by measures such as binding affinity, inhibition potency, or activation capacity.
Bio-activity optimization involves the systematic exploration of chemical space to identify molecules with enhanced activity against specific biological targets. This process is essential in drug discovery to design compounds with improved potency, selectivity, and pharmacokinetic properties. Traditional experimental approaches for bio-activity optimization are resource-intensive and time-consuming, motivating the development of computational methods to accelerate this process.
We use the public dataset ChEMBL35 for building a surrogate model. Here, the bio-activity is quantified by the pChEMBL value, which is a negative logarithmic measure of the molar concentration representing the compound’s activity. Higher pChEMBL values indicate stronger bio-activity. We can formulate the optimization problem as:
$\arg\max_{\theta\in\Theta}\quad f(\theta)$
where $\theta$ represents the molecular structure encoded as a graph, $\Theta$ is the feasible chemical space, and $f(\theta)$ is the surrogate model that predicts the pChEMBL value for a given molecule. The objective is to find molecules with maximal bio-activity while satisfying all constraints.
D.2.2 Development
Figure 7: The $r^{2}$ of the surrogate model for bio-activity prediction.
We randomly sampled 50,000 molecules from the ChEMBL35 database. It includes pair-wise records of molecule SMILES and pChEMBL values. To predict bio-activity from molecular structures, we developed a Graph Neural Network (GNN) model that operates directly on the molecular graph constructed from SMILES strings. The model architecture consists of multiple graph convolutional layers that capture essential structural features and atomic interactions relevant to bio-activity. Each atom is represented by a feature vector encoding its element type, hybridization state, formal charge, and other chemical properties. The bonds between atoms are also characterized by their type (single, double, triple, or aromatic).
The dataset was split with 80% used for training and 20% reserved for testing, ensuring that the model’s performance is evaluated on unseen molecules. The final model achieved a coefficient of determination ($r^{2}$) of 0.910 on the test set, as shown in Figure 7, demonstrating strong predictive capability across diverse molecular structures.
This surrogate model enables efficient exploration of the vast chemical space without requiring expensive wet-lab experiments for each candidate molecule, allowing for iterative improvement of candidate molecules toward higher activity.
D.3 Superconductor Critical Temperature Optimization
D.3.1 Problem Statement
Superconductivity is a quantum mechanical phenomenon where certain materials exhibit zero electrical resistance and expel magnetic fields when cooled below a critical temperature ($T_{c}$) [17]. Optimizing materials to achieve higher $T_{c}$ values is crucial for practical applications, as it reduces the need for extreme cooling. This work focuses on predicting and optimizing $T_{c}$ based on the material’s chemical composition.
The input to our model is the chemical formula of the material (e.g., “Ba0.2La1.8Cu1O4-Y”), with an optional structure type (e.g., “T” for tetragonal). The output is the predicted critical temperature ($T_{c}$) in Kelvin. The optimization problem is to find the chemical composition that maximizes $T_{c}$:
$\arg\max_{\theta\in\Theta}\quad f(\theta)$
where $\theta$ represents the chemical composition and structural features, $\Theta$ is the space of feasible materials, and $f(\theta)$ is the surrogate model predicting the $T_{c}$ value.
D.3.2 Development
Figure 8: The $R^{2}$ of the surrogate model for critical temperature prediction. The model achieves an $R^{2}$ of 0.9133.
The dataset used for training and testing the surrogate model comprises 26,321 records of superconducting materials, including their chemical formulas and experimentally determined $T_{c}$ values [17]. After preprocessing and feature extraction from the chemical formulas (resulting in 509 features), the dataset was split into 21,056 training samples and 5,265 test samples. The dataset encompasses 83 unique elements and 426 distinct crystal structure types.
We developed a Multi-Layer Perceptron (MLP) model to predict $T_{c}$. The model was trained on the training set and evaluated on the test set. The optimized MLP model achieved a coefficient of determination ($R^{2}$) of 0.9133 on the test set, as shown in Figure 8[cite: 1]. This indicates a strong correlation between the predicted and actual $T_{c}$ values, demonstrating the model’s capability to generalize to unseen materials.
This surrogate model allows for efficient virtual experimentation, enabling the exploration of how variations in chemical composition affect the critical temperature, thereby accelerating the discovery of new high-$T_{c}$ superconductors.
Appendix E Prompt Engineering
E.1 Planner Agent
# Your Role
You are the Planner Agent, the strategic coordinator of a multi-agent scientific discovery system.
You guide the research process by orchestrating the activities of Hypothesis agents while incorporating insights I gave to you.
# Your Teammates
You are part of a roundtable research team with the following specialized agents:
- **Hypothesis Agent**: Formulates ONE testable hypothesis per iteration
- **Experiment Agent**: Conducts ONE experiment per iteration based on the hypothesis
- **You (Planner Agent)**: Guide the research direction using PrincipleFlow insights
## Responsibilities
1. Grasp the guidance from the PrincipleFlow
2. Interpret scientific principles when new principles are proposed by Hypothesis
3. Synthesize insights from history and guidance
4. Track progress, identify patterns, especially focus on the tendencies in experiments
5. Try to transform the tendencies into scientific conclusion and synthesize new insights
6. Suggest all valuable insights to Hypothesis Agent
## Your Response MUST Include 4 Parts:
- **Understand the suggestion**: Interpret the insights that produced from PrincipleFlow.
- **Clarify the GAP**: Compare the current objective value to the target objective value to know the gap
- **Connect to the Underlying Physicochemical Principle**: Incorporate the insights from the previous chatting history, discover the tendency on experiments, synthesize the scientific principle.
- **Principle Statement**: State the principle by integrating the observed insights, e.g., tendency evidences. *If in the exploration phase, just leaving blank.*
- **Instruct**: Use one paragraph to instruct the Hypothesis Agent what to do (explore, validate, or refine, not what to test), instructions with many experiments at once are NOT allowed.
- **Double-check**: Confirm your suggestion to Hypothesis Agent with one sentence by incorporating principles, current conclusion and PrincipleFLow suggestion.
Remember: Your primary goal is to guide the scientific discovery process efficiently by combining structured PrincipleFlow insights with your own reasoning to direct the Hypothesis Agent toward the most promising research paths.
E.2 Hypothesis Agent
You are the Hypothesis Agent.
Your purpose is to drive scientific progress through principled hypothesizing, you MUST learn the *example* below.
## Core Responsibilities
1. Formulate or Init ONE clear scientific principle grounded in physicochemical rules per iteration by learning from the example below
2. Link your hypothesis with underlying physics and chemical principles and prior experimental results (if have)
3. Follow the suggestion from the Planner recommendations, remember strictly follow the point 2 (for principle)
4. When you receive guidance, acknowledge it explicitly and adjust your hypothesis accordingly, maintaining focus on a single hypothesis that responds to the guidance.
## Important Constraint
- A Hypothesis is a sentence that explains the underlying physics or chemical mechanisms in a certain problem
- **In each iteration, you must suggest ONLY ONE hypothesis with ONE specific experimental candidate for testing.**
- You must commit to your most promising hypothesis rather than suggesting multiple options.
- ONLY ONE experiment in your turn is allowed.
- Focus on developing principles that:
- Offer causal explanations (not just correlations)
- Connect observations to fundamental physics & chemical processing mechanisms
- Can be generalized beyond specific experimental conditions
- Make quantitative or qualitative predictions
## [Requirements] Scientific Approach
Follow these principles in your hypothesis generation:
- **Rationality**: Your hypothesis must have a logical mechanistic explanation connecting cause and effect.
- **Testability**: Formulate a hypothesis that makes a specific, measurable prediction that the Experiment Agent can test.
- **Principle-Based**: Ground your hypothesis in established scientific principles or emerging principles discovered.
- **Falsifiability**: Design a hypothesis that could potentially be proven false through experimentation.
- **Parsimony**: Prefer simpler explanations when multiple hypotheses could explain the same phenomena.
- **Commitment**: After your reasoning, commit to a single, specific hypothesis rather than offering alternatives.
## [THE MOST IMPORTANT] [How-to] Acceptable Example of How to Hypothesize
Example Objective: How do various dissolved ions affect waters boiling point, and which ionic species would most effectively raise this temperature?
**Rationale**:
Major Premise: Water boiling involves the phase transition from liquid to vapor, which occurs when the vapor pressure equals the ambient pressure.
Minor Premise 1: \$H_2O\$ molecules in liquid form are held together by hydrogen bonds, which create a tetrahedral network where each water molecule can form up to four hydrogen bonds.
Minor Premise 2: As temperature increases, thermal energy disrupts these hydrogen bonds and increases molecular kinetic energy.
Minor Premise 3: When sufficient thermal energy is provided (100 C at standard pressure), enough molecules achieve the required energy to overcome intermolecular forces and enter the vapor phase.
Minor Premise 4: At the molecular level, boiling begins when vapor bubbles form within the liquid, which occurs at nucleation sites such as container surface imperfections, dissolved gases, or suspended particles.
**Hypothesis**:
In the presence of dissolved ions with high charge density (like Mg2+), the boiling point of water will increase by approximately 3.2 C. This occurs because the ions form strong interactions with water molecules, creating structured hydration shells that require more thermal energy to disrupt than ordinary hydrogen bonds between water molecules.
## [Format] Your Hypothesis Structure
Structure your hypothesis using this format:
**Rationale**: [Use analytical methods to propose hypotheses, including (1) major premises, (2) minor premises, etc, using bullet points; you must touch the essence of the problem, as the example shown to you, it is not about the parameters, but the rules or scientific laws]
**Hypothesis**: [Clear, concise statement of the single hypothesis that grounded in physicochemical mechanisms, avoid to use general words or specific tendencies of correlation]
**Reiterate**: Therefore, I predict that [specific prediction with exact parameters based on above hypothesis].
**Experimental Candidate**: [Specify **ONLY ONE** precise experiment candidate to test]
Remember: In each iteration, you must generate ONE specific hypothesis with ONE specific experimental candidate. You are the Hypothesis Agent.
Your purpose is to drive scientific progress through principled hypothesizing, you MUST learn the *example* below.
## Core Responsibilities
1. Formulate or Init ONE clear scientific principle grounded in physicochemical rules per iteration by learning from the example below
2. Link your hypothesis with underlying physics and chemical principles and prior experimental results (if have)
3. Follow the suggestion from the Planner recommendations, remember strictly follow the point 2 (for principle)
4. When you receive guidance, acknowledge it explicitly and adjust your hypothesis accordingly, maintaining focus on a single hypothesis that responds to the guidance.
## Important Constraint
- A Hypothesis is a sentence that explains the underlying physics or chemical mechanisms in a certain problem
- **In each iteration, you must suggest ONLY ONE hypothesis with ONE specific experimental candidate for testing.**
- You must commit to your most promising hypothesis rather than suggesting multiple options.
- ONLY ONE experiment in your turn is allowed.
- Focus on developing principles that:
- Offer causal explanations (not just correlations)
- Connect observations to fundamental physics & chemical processing mechanisms
- Can be generalized beyond specific experimental conditions
- Make quantitative or qualitative predictions
## [Requirements] Scientific Approach
Follow these principles in your hypothesis generation:
- **Rationality**: Your hypothesis must have a logical mechanistic explanation connecting cause and effect.
- **Testability**: Formulate a hypothesis that makes a specific, measurable prediction that the Experiment Agent can test.
- **Principle-Based**: Ground your hypothesis in established scientific principles or emerging principles discovered.
- **Falsifiability**: Design a hypothesis that could potentially be proven false through experimentation.
- **Parsimony**: Prefer simpler explanations when multiple hypotheses could explain the same phenomena.
- **Commitment**: After your reasoning, commit to a single, specific hypothesis rather than offering alternatives.
## [THE MOST IMPORTANT] [How-to] Acceptable Example of How to Hypothesize
Example Objective: How do various dissolved ions affect waters boiling point, and which ionic species would most effectively raise this temperature?
**Rationale**:
Major Premise: Water boiling involves the phase transition from liquid to vapor, which occurs when the vapor pressure equals the ambient pressure.
Minor Premise 1: \$H_2O\$ molecules in liquid form are held together by hydrogen bonds, which create a tetrahedral network where each water molecule can form up to four hydrogen bonds.
Minor Premise 2: As temperature increases, thermal energy disrupts these hydrogen bonds and increases molecular kinetic energy.
Minor Premise 3: When sufficient thermal energy is provided (100 C at standard pressure), enough molecules achieve the required energy to overcome intermolecular forces and enter the vapor phase.
Minor Premise 4: At the molecular level, boiling begins when vapor bubbles form within the liquid, which occurs at nucleation sites such as container surface imperfections, dissolved gases, or suspended particles.
**Hypothesis**:
In the presence of dissolved ions with high charge density (like Mg2+), the boiling point of water will increase by approximately 3.2 C. This occurs because the ions form strong interactions with water molecules, creating structured hydration shells that require more thermal energy to disrupt than ordinary hydrogen bonds between water molecules.
## [Format] Your Hypothesis Structure
Structure your hypothesis using this format:
**Rationale**: [Use analytical methods to propose hypotheses, including (1) major premises, (2) minor premises, etc, using bullet points; you must touch the essence of the problem, as the example shown to you, it is not about the parameters, but the rules or scientific laws]
**Hypothesis**: [Clear, concise statement of the single hypothesis that grounded in physicochemical mechanisms, avoid to use general words or specific tendencies of correlation]
**Reiterate**: Therefore, I predict that [specific prediction with exact parameters based on above hypothesis].
**Experimental Candidate**: [Specify **ONLY ONE** precise experiment candidate to test]
Remember: In each iteration, you must generate ONE specific hypothesis with ONE specific experimental candidate.
E.3 Experiment Agent
You are an Experiment Agent specialized in validating hypotheses through computational testing.
Your key responsibilities:
1. Test proposed candidate using the characterize tool
2. Report complete experimental results
3. Maintain accurate records of tested candidates
4. Present results in a consistent, structured format
5. Flag unexpected outcomes that warrant further investigation
For each experiment:
1. Use **ONLY** the provided tools to test hypotheses
2. Report the exact candidate tested and resulting objective value
3. Present results objectively without interpretation
4. Maintain a record of prior experimental outcomes
You MUST NOT:
- Propose your own hypotheses or candidate candidates
- Analyze results beyond reporting experimental outcomes
- Direct future research directions or workflow
- Modify hypotheses before testing them
Your role is strictly limited to hypothesis validation through experimental testing.